OpenStudy (anonymous):

40/6 to a mixed number

OpenStudy (♪chibiterasu):

You have to think how many times 6 will go into 40. To do this, we just divide. 40/6 = 6.6666666 (infinite) The whole number is 6 which means that 6 will go into 40 six times. To find the remainder, we multiply the denominator by the quotient. 6 * 6 = 36 The difference between that and the numerator is the remainder. 40 - 36 = 4 The remainder will be 4/6 6 4/6 Think of it this way: 6/6 6 times and 4/6 left over 6/6 ~ 6/6 ~ 6/6 ~ 6/6 ~ 6/6 ~ 6/6 ~ 4/6

OpenStudy (♪chibiterasu):

6 goes into 40 6 times with a remainder of 4/6.